# Boundary layers in thermal convection are fluctuation-dominated

### v1

@@ Abstract
We study the dynamics of thermal and momentum boundary layers in three-dimensional direct numerical simulations of Rayleigh-Bénard convection for the Rayleigh number range $10^{5}\leq Ra\leq 10^{11}$ and $Pr=0.7$. Using a Cartesian slab with horizontal periodic boundary conditions and an aspect ratio of 4, we obtain statistical homogeneity in the horizontal $x$- and $y$-directions, thus approximating an infinitely extended system. We observe that upon canonical use of long-time and area averages, a coherent mean flow is practically absent. Instead, the velocity field close to the wall is a collection of differently oriented local shear patches interspersed by shear-free, incoherent flow regions. These shear patches occupy an area fraction of approximately 40% for all $Ra$. Rather than resulting in a pronounced mean with small fluctuations about it, the velocity field is dominated by strong fluctuations of all three components around a non-existent or weak mean. This feature is particularly pronounced for $Ra\geq 10^{9}$. We discuss the consequences of these observations for convection layers with larger aspect ratios, including boundary layer instabilities and the resulting turbulent heat transport.
I Introduction
@@
Rayleigh-Bénard convection (RBC) is one of the fundamental flow configurations in fluid turbulence research. The fluid in this configuration is nominally confined to an infinitely extended layer enclosed between two parallel, horizontal and impermeable plates separated by a vertical distance $H$ (Rayleigh, 1916). When the fluid layer is heated sufficiently strongly from below (and cooled from above), buoyancy forces initiate a turbulent fluid motion that has a statistically preferred state with respect to the direction of gravity, ${\bm{g}}={g\bm{e}}_{z}$. A central question concerns the amount of heat and momentum carried through the layer, and their dependencies on the imposed temperature difference between the top and bottom plates, $\Delta T=T_{\rm bot}-T_{\rm top}$. The temperature difference is expressed by the dimensionless Rayleigh number $Ra=g\alpha\Delta TH^{3}/(\nu\kappa)$, where $\alpha$ is the isobaric expansion coefficient, $\nu$ the kinematic viscosity, and $\kappa$ the temperature diffusivity (Ahlers et al., 2009a; Chillà and Schumacher, 2012; Verma, 2018).
@@
Since the RBC system is enclosed by walls at the top and bottom, the viscous and thermal boundary layers formed on these walls pose a bottleneck for the global transport of both heat and momentum. Their composition and dynamics at very high Rayleigh numbers still need to be better understood as emphasized recently (Iyer et al., 2020; Lindborg, 2023; Shishkina and Lohse, 2023; Creyssels and Martinard, 2024). For Rayleigh numbers $Ra\gtrsim 10^{10}$, no laboratory experiment to-date has resolved the dynamic interplay of the boundary layers and their fluctuations with the basic structural elements, namely thermal plumes and shear layers. Direct numerical simulations (DNS) are thus the only way to compare their structure and statistical properties with predictions from theories for canonical laminar and turbulent boundary layers (Schlichting and Gersten, 2016). Furthermore, the closed-cell geometry of high-Rayleigh-number studies (Castaing et al., 1989; Chavanne et al., 1997; Niemela et al., 2000; Ahlers et al., 2009b; Urban et al., 2012) breaks the horizontal translation symmetry of the statistics, except possibly when the aspect ratio is very large (Pandey et al., 2018). Small aspect ratio enforces a dominant large-scale circulation (LSC) in the cell (Kadanoff, 2001), manifesting as a relatively coherent shear flow connecting the top and bottom plates, fluctuating only moderately in its mean orientation (Sreenivasan et al., 2002; Stevens et al., 2011; Shi et al., 2012; Scheel and Schumacher, 2017). Both aspects take us away from the original question on the heat and momentum transfer in an infinitely extended plane layer.
In this work, we focus on a configuration that is closer to the original RBC model of convection between a pair of infinitely extended planes, by using periodic boundaries in horizontal directions. Simulations with similar boundary conditions and Prandtl numbers have been done by Kerr (1996) for $\Gamma=6$ and $Ra\leq 2\times 10^{7}$, by Hartlep et al. (2003) for $\Gamma=10$ and $Ra\leq 1\times 10^{7}$, by van Reeuwijk et al. (2008) and van Reeuwijk and Jonker (2008) for $\Gamma=4$ and $Ra\leq 1\times 10^{8}$, and by De et al. (2018) for $\Gamma=6$ but $Ra\leq 2\times 10^{6}$. Our DNS spans Rayleigh numbers of six orders of magnitude up to $Ra=10^{11}$ for long periods of time (see table 1 for details). A canonical mean flow analysis reveals practically no mean flow; instead, strong velocity fluctuations dominate the flow at all $Ra$. Fits to the mean velocity profiles result in very small free-stream velocities $U_{\infty}\sim 10^{-3}$ in terms of the free fall velocity $U_{f}=\sqrt{g\alpha\Delta TH}$, and thicknesses $\delta_{\infty}\sim 10^{-2}$ in terms of $H$, resulting in small shear Reynolds numbers $Re_{\rm shear}\lesssim 1$–10 even for the largest $Ra$. Furthermore, we analyse fluctuations of the velocity components, determine the distances of maximum mean-square fluctuation from the wall, and discuss the resulting Reynolds numbers. A first impression of the complex boundary layer dynamics can be obtained by the streamline plots in figure 1 at $z=\delta_{T}/2$ for two $Ra$. The figures indicate a prominent patchiness of the boundary layer viewed as a whole from the top. We will quantify the corresponding area fraction and relate it to the large-scale patterns, which are the turbulent superstructures of convection (Pandey et al., 2018).
II Numerical simulations
Figure 1: Streamline snapshots of the horizontal velocity field $(u_{x},u_{y})$ at $z=\delta_{T}/2$. (a) $Ra=10^{8}$. (b) $Ra=10^{11}$. The entire cross section of $4H\times 4H$ is shown.
@@ II Numerical simulations
We solve the three-dimensional Boussinesq equations of RBC (Verma, 2018) by the GPU-based spectral element code nekRS (Fischer et al., 2022). Length, velocity, and temperature are expressed in units of $H$ and $U_{f}$, and the outer temperature difference $\Delta T$. No-slip boundary conditions apply for the velocity field at the plates at $z=0$ and $H$. Table 1 summarizes 7 simulations, all at a Prandtl number $Pr=\nu/\kappa=0.7$ and aspect ratio $\Gamma=L/H=4$, where $L$ is the horizontal length. The number of collocation points inside the thermal boundary layer (based on the temperature fluctuation profiles) is always $N_{BL}\geq 11$. Furthermore, we verified that the Nusselt numbers $Nu_{\rm vol}$ and $Nu_{\rm wall}$, which are given by combined volume-time $\langle\cdot\rangle_{V,t}$ and area-time averages $\langle\cdot\rangle_{A,t}$,
@@
Figure 2(a,b) plot the Nusselt and Reynolds numbers, $Nu=Nu_{\rm wall}$ and $Re=U_{\rm rms}\sqrt{Ra/Pr}$ versus the Rayleigh number $Ra$, compensated by the high-$Ra$ scaling result of Iyer et al. (2020); here $U_{\rm rms}=\langle{\bm{u}}^{2}\rangle_{V,t}^{1/2}$. We also compare our results with those of Scheel and Schumacher (2017) in a closed cylinder at $\Gamma=1$. While the Nusselt numbers collapse fairly well (figure 2(a)) and are in agreement with previous simulations for similar parameters and those for $Ra\leq 10^{8}$ from van Reeuwijk et al. (2008), the Reynolds number shows a strong geometry dependence (figure 2(b)): they show good agreement in scaling exponents at high $Ra$ but not in the prefactor. Figure 2(c) shows the thermal fluctuation boundary layer thickness-based Rayleigh number $Ra_{\delta,{\rm rms}}=g\alpha(\Delta T/2)\delta^{3}_{T,{\rm rms}}/(\nu\kappa)$ versus $Ra$, together with the power law fit. Statistics in the all runs are obtained for more than 100 free-fall times $T_{f}=H/U_{f}$; see table 1. We also verified that the vertical profiles of the temperature fluctuation, fitted by polynomial orders $p=7$ and 9 at $Ra=10^{10}$ and $p=5,7$ and 9 at $Ra=10^{11}$, collapse on each other. These observations confirm that the DNS are sufficiently well-resolved.
@@
Figure 2: (a) Compensated turbulent heat transfer $Nu/Ra^{1/3}$ versus $Ra$. (b) Compensated turbulent momentum transfer $Re/Ra^{0.458}$ versus $Ra$. (c) Thermal fluctuation boundary layer thickness-based Rayleigh number $Ra_{\delta,{\rm rms}}$ versus $Ra$, together with a power law fit. Panels (a,b) contain data from the present DNS and from Scheel and Schumacher (2017) and Iyer et al. (2020), compensated by the fits to the data from Iyer et al. (2020); see also table 1.
$Ra$ $N_{e}$ $p$ $N_{BL}$ $\Delta t$ $Nu_{\mathrm{vol}}$ $Nu_{\mathrm{wall}}$ $Re$ $Ra_{\delta,{\rm rms}}$
$10^{5}$ $100\times 100\times 64$ 5 71 1000 $4.27\pm 0.24$ $4.27\pm 0.15$ $93.0\pm 2.6$ $102\pm 19$
$10^{6}$ $100\times 100\times 64$ 7 57 1000 $8.15\pm 0.35$ $8.15\pm 0.19$ $296\pm 5$ $115\pm 15$
$10^{7}$ $100\times 100\times 64$ 9 42 1000 $15.6\pm 0.9$ $15.6\pm 0.3$ $892\pm 17$ $144\pm 14$
$10^{8}$ $150\times 150\times 96$ 7 24 1000 $30.4\pm 1.9$ $30.4\pm 0.4$ $2571\pm 47$ $175\pm 15$
$10^{9}$ $150\times 150\times 96$ 9 16 400 $60.9\pm 4.4$ $60.8\pm 0.9$ $7300\pm 150$ $209\pm 15$
$10^{10}$ $250\times 250\times 128$ 9 13 200 $123.5\pm 9.7$ $123.1\pm 1.4$ $20510\pm 470$ $258\pm 36$
$10^{11}$ $500\times 500\times 256$ 5 11 100 $255\pm 19$ $256\pm 3$ $57960\pm 870$ $277\pm 22$
Table 1: Details of the simulation series. The Rayleigh number $Ra$, the number of spectral elements $N_{e}$, the polynomial order $p$ on each element with respect to the space direction, the number of collocation points within the thermal boundary layer $N_{BL}$, the total averaging time in free-fall units $\Delta t$, the volume averaged and the wall-averaged Nusselt numbers, $Nu$ and $Nu_{\mathrm{wall}}$, the Reynolds number $Re$, and the Rayleigh number based on the thermal fluctuation boundary layer thickness, $Ra_{\delta,{\rm rms}}$. The total number of mesh cells is $N_{e}\times p^{3}$. It increases to almost $2.2\times 10^{10}$ collocation mesh cells for $Ra=10^{11}$. Mean values in the last four columns are accompanied by the standard deviations.
@@
Figure 3: (a) Planar-averaged profiles of the horizontal velocity component $u_{x}$ for different averaging intervals for $Ra=10^{9}$. (b) Match of the profiles from panel (a) to the Blasius profile $f^{\prime}(\eta)$ (dashed line). Notice the very small values of $U_{\infty}$ and the non-monotonic convergence of the profiles. (c) Variation of the free-stream velocity $U_{\infty}$ and boundary layer height $\delta_{\infty}$ obtained from the fits to the Blasius profile versus averaging time. (d) Time-averaged profiles $\langle u_{x}\rangle_{A,t}$ for all 7 Rayleigh numbers. (e,f) Combined plane-time average profile (thick red line) and instantaneous plane average profiles (80 thin grey lines) of the horizontal velocity fluctuations, $U^{h}_{\rm rms}(z)$ and $U_{\rm rms}(z)$ for $Ra=10^{9}$.
III Mean temperature and velocity profiles

### v2

# No sustained mean velocity in the boundary region of plane thermal convection

@@ Abstract
We study the dynamics of thermal and momentum boundary regions in three-dimensional direct numerical simulations of Rayleigh-Bénard convection for the Rayleigh number range $10^{5}\leq Ra\leq 10^{11}$ and $Pr=0.7$. Using a Cartesian slab with horizontal periodic boundary conditions and an aspect ratio of 4, we obtain statistical homogeneity in the horizontal $x$- and $y$-directions, thus approximating best an extended convection layer relevant for most geo- and astrophysical flow applications. We observe upon canonical use of combined long-time and area averages, with averaging periods of at least 100 free-fall times, that a global coherent mean flow is practically absent and that the magnitude of the velocity fluctuations is larger than the mean by up to 2 orders of magnitude. The velocity field close to the wall is a collection of differently oriented local shear-dominated flow patches interspersed by extensive shear-free incoherent regions which can be as large as the whole cross section, unlike for a closed cylindrical convection cell of aspect ratio of the order 1. The incoherent regions occupy a 60% area fraction for all Rayleigh numbers investigated here. Rather than resulting in a pronounced mean flow with small fluctuations about such a mean, as found in small-aspect-ratio convection, the velocity field is dominated by strong fluctuations of all three components around a non-existent or weak mean. We discuss the consequences of these observations for convection layers with larger aspect ratios, including boundary layer instabilities and the resulting turbulent heat transport.
keywords:
Rayleigh-Bénard convection, boundary layers
1 Introduction
@@
Rayleigh-Bénard convection (RBC) is one of the fundamental flow configurations in fluid turbulence research. The fluid in this configuration is nominally confined to an infinitely extended layer enclosed between two parallel, horizontal and impermeable plates separated by a vertical distance $H$ (Rayleigh, 1916). When the fluid layer is heated sufficiently strongly from below (and cooled from above), buoyancy forces initiate a turbulent fluid motion that has a statistically preferred state with respect to the direction of gravity, ${\bm{g}}={g\bm{e}}_{z}$. A central question concerns the amount and nature of heat and momentum carried through the layer, and their dependencies on the imposed temperature difference between the top and bottom plates, $\Delta T=T_{\rm bot}-T_{\rm top}$. The temperature difference is expressed by the dimensionless Rayleigh number $Ra=g\alpha\Delta TH^{3}/(\nu\kappa)$, where $\alpha$ is the isobaric expansion coefficient, $\nu$ the kinematic viscosity, and $\kappa$ the temperature diffusivity (Ahlers et al., 2009b; Chillà & Schumacher, 2012; Verma, 2018).
@@
A second important question is the structure of the velocity and thermal boundary layers on the horizontal walls and their effects on heat transport. Since the RBC system is enclosed by walls at the top and bottom, the viscous and thermal boundary layers formed on these walls pose a bottleneck for the global transport of both heat and momentum. Their composition and dynamics at very high Rayleigh numbers still need to be better understood as emphasized recently (Iyer et al., 2020; Lindborg, 2023; Shishkina & Lohse, 2023; Creyssels & Martinard, 2024). For Rayleigh numbers $Ra\gtrsim 10^{10}$, no laboratory experiment to-date has resolved the dynamic interplay of the boundary layers and their fluctuations with the basic structural elements, namely thermal plumes and shear layers. Direct numerical simulations (DNS) are thus the only way to compare their structure and statistical properties with predictions from theories for canonical laminar and turbulent boundary layers (Schlichting & Gersten, 2016). Furthermore, the closed-cell geometry of high-Rayleigh-number studies (Castaing et al., 1989; Chavanne et al., 1997; Niemela et al., 2000; Ahlers et al., 2009a; Urban et al., 2012) breaks the horizontal translation symmetry of the statistics, except possibly when the aspect ratio is very large (Pandey et al., 2018). Small aspect ratio enforces a dominant large-scale circulation (LSC) in the cell (Kadanoff, 2001), manifesting as a relatively coherent shear flow connecting the top and bottom plates, fluctuating only moderately in its mean orientation (Sreenivasan et al., 2002; Stevens et al., 2011; Shi et al., 2012; Scheel & Schumacher, 2017). Both aspects take us away from the original question on the heat and momentum transfer in an infinitely extended plane layer.
In this work, we focus more on the second question in a configuration that is closest to the original RBC model of convection between a pair of infinitely extended planes, by using periodic boundaries in both horizontal directions. Simulations with similar boundary conditions and Prandtl numbers have been done by Kerr (1996) for $\Gamma=6$ and $Ra\leq 2\times 10^{7}$, by Hartlep et al. (2003) for $\Gamma=10$ and $Ra\leq 1\times 10^{7}$, by van Reeuwijk et al. (2008a) and van Reeuwijk et al. (2008b) for $\Gamma=4$ and $Ra\leq 1\times 10^{8}$, by De et al. (2018) for $\Gamma=6$ but $Ra\leq 2\times 10^{6}$, and by Stevens et al. (2018) for $\Gamma\leq 32$ at $Ra=10^{9}$. Our DNS spans Rayleigh numbers of six orders of magnitude up to $Ra=10^{11}$ for long periods of time (see table 1 for details). The choice of an aspect ratio of 4 for the present study provides a “sweet spot”. On the one hand, the domain is large enough that it does not generate strong large scale circulations, see Niemela & Sreenivasan (2006). On the other hand, it is small enough to allow us to advance to very high Rayleigh numbers, here of up to $Ra=10^{11}$, since the required numerical resources grow with $\Gamma^{2}$. Furthermore, this is the aspect ratio beyond which Nusselt and Reynolds numbers become independent of $\Gamma$ as discussed in Stevens et al. (2018). We supplement these results by additional DNS runs at $\Gamma=2$ and 8 for $Ra=10^{9}$.
A canonical mean flow analysis reveals practically no global mean flow; instead, strong velocity fluctuations dominate the flow at all $Ra$. Fits to the mean vertical velocity profiles result in very small free-stream velocities $U_{\infty}\sim 10^{-3}$ in terms of the free fall velocity $U_{f}=\sqrt{g\alpha\Delta TH}$, and thicknesses $\delta_{\infty}\sim 10^{-2}$ in terms of $H$, resulting in small shear Reynolds numbers $Re_{\rm shear}\lesssim 1$–10 even for the largest $Ra$. We further analyse fluctuations of the velocity components, determine the distances of maximum mean-square fluctuations from the wall, and discuss the resulting Reynolds numbers.
Figure 1: Turbulent fields at two Rayleigh numbers. Streamline snapshots of the horizontal velocity field $(u_{x},u_{y})$ and contour plots of temperature field $T$ at the same instant. (a,c) $Ra=10^{8}$. (b,d) $Ra=10^{11}$. The entire cross section of $4H\times 4H$ at $z=\delta_{T}/2$ close to the bottom wall is shown. $H$ is the height of the convection layer and $\delta_{T}$ the thermal boundary layer thickness. The colour legend holds for both temperature field panels (c,d). In panels (a) and (b) examples for local coherent shear-dominated and incoherent shear-free regions are indicated.
A quick impression of the complex boundary layer dynamics is obtained by the streamline and contour plots in figure 1 close to the bottom wall for two Rayleigh numbers. The figures indicate a prominent patchiness of the whole velocity boundary layer viewed from the top. The boundary layer is composed of coherent shear-dominated and incoherent shear-free regions. This feature becomes less prominent for the contours of the temperature field, which display an increasingly dense skeleton of thermal plume ridges over the whole plate. Here, we quantify the corresponding area fraction, condition the fluctuations on the coherent and incoherent regions, and relate the incoherent regions to the large-scale patterns in the bulk, which are the turbulent superstructures of convection (Pandey et al., 2018; Stevens et al., 2018). Only when the velocity fluctuations are conditioned on the coherent shear-dominated regions are the mean profiles close to those observed in closed cylinders for $\Gamma\sim 1$ (Scheel & Schumacher (2017).
One important point needs to be made here. While a thermal boundary layer of the standard type is indeed present, no momentum boundary layer can be easily identified, as explained subsequently. For the velocity field, it is thus more accurate to merely discuss the flow near the wall instead of the boundary layer flow, but for convenience of identification and following convention, we continue to use the term boundary layer. There is no such ambiguity for the thermal boundary layer.
2 Numerical simulations and resolution analysis
We solve the three-dimensional Boussinesq equations of RBC (Verma, 2018) by the GPU-based spectral element code nekRS (Fischer et al., 2022) which combines an element decomposition of the computational domain with a spectral expansion in Lagrange polynomials of each involved field along each spatial dimension on each element. The equations are given in dimensionless form by
$\frac{\partial{\bm{u}}}{\partial t}+({\bm{u}}\cdot{\bm{\nabla}}){\bm{u}}$ $=-{\bm{\nabla}}p+T\hat{{\bm{z}}}+\sqrt{\frac{Pr}{Ra}}\,\nabla^{2}{\bm{u}},$ (1)
$\frac{\partial T}{\partial t}+({\bm{u}}\cdot{\bm{\nabla}})T$ $=\frac{1}{\sqrt{PrRa}}\,\nabla^{2}T,$ (2)
${\bm{\nabla}}\cdot{\bm{u}}$ $=0.$ (3)
@@ 2 Numerical simulations and resolution analysis
Here, ${\bm{u}}$, $p$, and $T$ are the velocity, pressure, and temperature fields, respectively. Length, velocity, and temperature are expressed in units of $H$, $U_{f}$, and the outer temperature difference $\Delta T$, respectively. No-slip boundary conditions apply for the velocity field at the plates at $z=0$ and $H$. Table 1 summarizes 10 simulations, all at a Prandtl number $Pr=\nu/\kappa=0.7$ and aspect ratios $\Gamma=L/H=2$, 4, and 8 where $L$ is the horizontal length. The number of collocation points inside the thermal boundary layer (based on the temperature fluctuation profiles) is always $N_{BL}\geq 15$. Furthermore, we verified that the Nusselt numbers $Nu_{\rm vol}$ and $Nu_{\rm wall}$, which are given by combined volume-time $\langle\cdot\rangle_{V,t}$ and area-time averages $\langle\cdot\rangle_{A,t}$,
@@
As a first global result, we plot in figure 2(a,b) the Nusselt and Reynolds numbers, $Nu=Nu_{\rm vol}$ and $Re=U_{\rm rms}\sqrt{Ra/Pr}$ versus the Rayleigh number $Ra$, compensated by the high-$Ra$ scaling result of Iyer et al. (2020); here $U_{\rm rms}=\langle{\bm{u}}^{2}\rangle_{V,t}^{1/2}$. Also compared are our results with those of Stevens et al. (2010) and Scheel & Schumacher (2017) in closed cylinders at $\Gamma=1/2$ and $\Gamma=1$, respectively. Considering the large differences in aspect ratios, the Nusselt numbers collapse fairly well and follow the same trend for $Ra\leq 10^{11}$ (figure 2(a)); they are also in agreement with previous simulations for $Ra\leq 10^{8}$ from van Reeuwijk et al. (2008a). The Reynolds number shows a strong geometry dependence as visible in figure 2(b), though they tend to the same exponent towards high $Ra$; this suggests an agreement in scaling exponents at high $Ra$, but the prefactor seems to have a complicated Rayleigh number dependence. Statistics in the all runs are obtained for equal and more than 100 free-fall times $T_{f}=H/U_{f}$; see table 1.
We verified that the resolution of the boundary layers is sufficient. Figure 3 shows the vertical profiles of the temperature fluctuations for three different spectral element grids at $Ra=10^{10}$ and two at $Ra=10^{11}$ with different polynomial orders $p\geq 5$. It is seen that the profiles collapse well on each other, thus also demonstrating the convergence of the spectral method. This conclusion is further reinforced for $Ra=10^{9}$ where we have 4 DNS runs at different aspect ratio and vertical grid stretching. First, the Nusselt and Reynolds numbers in table 1 differ only slightly (of the order of a tenth of a percent) between $\Gamma=4$ and 8. Secondly, they are very close for two element grids with different vertical grid stretching at $\Gamma=4$. The fast spectral convergence of spectral element methods in comparison to lower-order finite-difference schemes has been reported recently by Zahtila et al. (2023) in comprehensive studies for turbulent channel flows.
@@
Figure 2: (a) Compensated turbulent heat transfer $Nu/Ra^{0.331}$ versus $Ra$. (b) Compensated turbulent momentum transfer $Re/Ra^{0.458}$ versus $Ra$. The panels contain data from the present DNS and from Scheel & Schumacher (2017), Stevens et al. (2010) and Iyer et al. (2020). All data are compensated by the fits to the data from Iyer et al. (2020); see also table 1.
Figure 3: Resolution study for the runs at the two highest Rayleigh numbers. The near-wall profile of the temperature fluctuations is plotted. The legend indicates the Rayleigh number together with the vertical numbers of spectral elements and the polynomial order. The runs, as listed in the legend from top to bottom, have 11, 14, 16, 11, and 16 points inside the thermal boundary layer. The horizontal lines are the thermal fluctuation boundary layer thicknesses.
[…]
$10^{5}$ 4 $100\times 100\times 64$ 5 71 1000 $4.27\pm 0.24$ $4.27\pm 0.15$ $93.0\pm 2.6$ $104\pm 21$
$10^{6}$ 4 $100\times 100\times 64$ 7 57 1000 $8.15\pm 0.35$ $8.15\pm 0.19$ $296\pm 5$ $116\pm 14$
$10^{7}$ 4 $100\times 100\times 64$ 9 42 1000 $15.6\pm 0.9$ $15.6\pm 0.3$ $892\pm 17$ $144\pm 14$
$10^{8}$ 4 $150\times 150\times 96$ 7 24 1000 $30.4\pm 1.9$ $30.4\pm 0.4$ $2571\pm 47$ $174\pm 15$
$10^{9}$ 4 $150\times 150\times 96$ 9 15 400 $60.9\pm 4.4$ $60.8\pm 0.9$ $7300\pm 150$ $209\pm 15$
$10^{9\,\dagger}$ 2 $120\times 120\times 96$ 9 21 100 $62.4\pm 5.0$ $62.4\pm 1.2$ $6990\pm 210$ $177\pm 21$
$10^{9\,\dagger}$ 4 $220\times 220\times 96$ 9 22 100 $60.5\pm 3.8$ $60.5\pm 0.7$ $7340\pm 140$ $206\pm 18$
$10^{9\,\dagger}$ 8 $450\times 450\times 96$ 9 22 100 $60.9\pm 2.0$ $60.9\pm 0.3$ $7180\pm 70$ $212\pm 10$
$10^{10}$ 4 $400\times 400\times 200$ 7 16 100 $122.9\pm 10.8$ $123.1\pm 1.4$ $20720\pm 380$ $247\pm 22$
$10^{11}$ 4 $500\times 500\times 256$ 7 16 100 $254\pm 21$ $254\pm 2$ $58550\pm 1040$ $298\pm 23$
[…]
@@
Figure 4: (a) Planar-averaged profiles of the horizontal velocity component $u_{x}$ for different averaging intervals for $Ra=10^{9}$. Notice the very small values of $U_{\infty}$ and the non-monotonic convergence of the profiles. (b) Match of the near-wall profiles from panel (a) to the Blasius profile $f^{\prime}(\eta)$ (dashed line). (c) Variation of the free-stream velocity $U_{\infty}$ and boundary layer height $\delta_{\infty}$ obtained from the fits to the Blasius profile versus averaging time at $Ra=10^{9}$. (d) Time-averaged profiles $\langle u_{x}\rangle_{A,t}$ for all 7 Rayleigh numbers. (e) Time-averaged profiles $\langle u_{y}\rangle_{A,t}$ for all 7 Rayleigh numbers. (f) Time-averaged profiles $\langle u_{x}\rangle_{A,t}$ for Rayleigh number $Ra=10^{9}$ and aspect ratios $\Gamma=2$, 4 and 8. These are the three runs which are indicated by a dagger in table 1.
3 Mean profiles of temperature and velocity